Friendly’s Quick Loans, Inc., offers you “thirteen for sixteen or I knock on your door.” This means you get $13 today and repay $16 when you get your paycheck in one week (or else).

If you were brave enough to ask, what APR would Friendly’s say you were paying? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

APR            %

What’s the effective annual return Friendly’s earns on this lending business? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

EAR            %
Answer #1

1) APR means annual percentage return, if we calculate weekly return, we can calculate yearly return by multiplying the weekly return by 52 weeks,

to calculate 1-week return, we can use the RATE formula in excel,

=RATE(time,pmt,present value,future value)

=RATE(1,0,-13,16)

=23.07692%

APR = 23.07692% * 52 weeks = 1200%

2)

EAR = (1+r)^52 - 1

=(1+0.230769231)^52 -1

=4888503.09%